OpenStudy (anonymous):

What is the slope of the line that passes through the points (−5,4) and (15,−4)? A. -2/5 B. 0 C. -5/2 D. undefined

OpenStudy (anonymous):

Well, let's start with this formula to get the slope:\[ m=\frac{y2-y1}{x2-x1}\]

sam (.sam.):

\[\huge Slope=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]

OpenStudy (anonymous):

(-5, 4) & (15,-4) We could say that x1 = -5 y1 = 4 x2 = 15 y2 = -4

OpenStudy (anonymous):

So from there: \[ m=\frac{y2-y1}{x2-x1}\] \[ m=\frac{(-4)-(4)}{(15)-(-5)}\] \[ m=\frac{-4-4}{15+5}\] \[ m=\frac{-8}{20}\] \[ m=-\frac{2}{5}\]

OpenStudy (anonymous):

Does that make sense?